OpenStudy (konradzuse):

How to solve for sin(x)/(cos^2(x))?

OpenStudy (konradzuse):

apparently it's sec(x)tan(x). I understand that 1/cos(x) = sec(x) and 1/cos^2(x) = sec(x) but why with the sin(x)/cos(x) do we get 1 sec and 1 tan...?

OpenStudy (phi):

isn't it just \[ \frac{sin(x)}{cos(x)}\cdot \frac{1}{cos(x)}\]

OpenStudy (konradzuse):

is there a sign in the middle of that?

OpenStudy (phi):

no, multiplying

OpenStudy (konradzuse):

I was thinkikng that actually, thanks.

OpenStudy (konradzuse):

Now that I think about it, I just realized it's just sec(x)tan(x)... I thought it was sec^2(x) tan(x) nvm :p

OpenStudy (konradzuse):

derpa thanks Phi :P.

OpenStudy (phi):

yes sin/cos = tan and 1/cos = sec

OpenStudy (konradzuse):

mhmhm :)

OpenStudy (konradzuse):

Actually, I was looking at wolfram and the answer key and realized I was konfused...

OpenStudy (konradzuse):

wolfram said sin(x)/cos^(x) is = sec(x)tan(x).

OpenStudy (konradzuse):

the actual question was (1)/(1-sin(x))

OpenStudy (konradzuse):

it got to a point where it was int of 1/cos^2(x) + int sin(x)/cos^2(x)

OpenStudy (konradzuse):

the professor changed it to - int -sin(x)/cos^2(x)

OpenStudy (konradzuse):

then did -cos^-1(x)/-1?

OpenStudy (konradzuse):

which turned into 1/cos(x) which turned into sec(x)....

OpenStudy (konradzuse):

@phi what did she do?

OpenStudy (phi):

exactly what was (is) the original question? is it an integral?

OpenStudy (konradzuse):

KonradZuse 0 the actual question was int (1)/(1-sin(x))

OpenStudy (konradzuse):

1/(1-sin(x)) * (1+sin(x))/(1+sin(x))

OpenStudy (konradzuse):

(1+sin(x))/(1-sin^2(x))

OpenStudy (konradzuse):

(1+sin(x))/cos^(x)

OpenStudy (konradzuse):

(1)/cos^2(x) + (sin(x))/(cos^2(x))

OpenStudy (konradzuse):

first part is sec^2(x) second part would be sec(x)tan(x)... Problem is you need to integrate it...

OpenStudy (konradzuse):

:(?

OpenStudy (phi):

I would write sec(x)tan(x) as \[\int cos^{-2}(x) sin(x) dx \] now this is in the form \[ \int u^{-2} du \]

OpenStudy (phi):

with a minus sign in there

OpenStudy (phi):

u = cos(x) du = -sin(x) dx

OpenStudy (konradzuse):

oic that makes snese now, thanks :)